what is the surface area of a cell if the diameter is 30 and time to center is 0.075?
Answer #1

Considering cell to be spherical, its surface area (S) can be calculated using the formula,

S=4πr^2

r=diameter/2

r=30/2 =15

S=4×(22/7) ×(15×15)

S = 2828.57 square units.